when pendulum is vibrating then no sound is heard why ?
1
Expert's answer
2019-07-15T09:12:44-0400

Human hearing range is from 20 to 20000 Hz, although there is variation between individuals.

Let's consider for example pendulum with the 5 cm length. The frequency of oscillations of such pendulum is

"F=\frac{1}{2 \pi \sqrt{\frac{l}{g}}}= \frac{1}{2 \pi \sqrt{\frac{0.05 m}{9.8 m/s^2}}} \approx 2.2 Hz"

So we can see that the frequency is too low to we can hear the sound.